OpenStudy (anonymous):

The center of the fountain is located at (-4,8). The radius of the fountain is 25 feet. What equation could represent the edge of the fountain ? A. (x+4)^2+(y-8)^2=5 B. (x+4)^2+(y-8)^2=25 C. (x+4)^2+(y-8)^2=625 D. (x-4)^2+(y+8)^2=625

OpenStudy (campbell_st):

well the general form of a circle with centre at (h, k) and radius r is \[(x - h)^2 + (y - k)^2 = r^2\] substitute the information you have been given and then simplify

OpenStudy (imstuck):

I don't even know that you have to know the center of the circle as long as you have that the radius is 25. Using the circumference formula, we get C = pi(50) for around that circle with a radius of 25. I can't see any other way to do this, unless I am missing something somewhere!